An ideal gas is 45 psig and 80°F is heated in a closed container to 130°F. What is the final pressure in psia?


1
Expert's answer
2021-06-07T15:01:42-0400
"\frac{p_1}{T_1}=\frac{p_2}{T_2}\\\frac{45}{80+ 459.67}=\frac{p_2}{130+ 459.67}\\p_2=49\ psig"